OpenStudy (yanasidlinskiy):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd?

OpenStudy (yanasidlinskiy):

@jim_thompson5910 can you help??

jimthompson5910 (jim_thompson5910):

What is the probability Jason picks an odd number?

OpenStudy (yanasidlinskiy):

Umm...I don't know....

jimthompson5910 (jim_thompson5910):

which numbers in the list below are odd 1,2,3,4,5,6,7,8,9,10

OpenStudy (yanasidlinskiy):

1,3,5,7,9

OpenStudy (yanasidlinskiy):

So it would be 5

jimthompson5910 (jim_thompson5910):

so you listed 5 numbers, that's out of 10 total the probability Jason picks an odd number is 5/10 = 1/2

OpenStudy (yanasidlinskiy):

Wow.....what an easy question I posted........So dumb.....Thanx!!!!

jimthompson5910 (jim_thompson5910):

1/2 isn't the final answer though (it's going to get us there)

jimthompson5910 (jim_thompson5910):

you do about the same steps to find the probability of Kyle picking an odd number to also be 1/2

OpenStudy (yanasidlinskiy):

So 1/4 right?

jimthompson5910 (jim_thompson5910):

so the probability of them both picking an odd number is (1/2) * (1/2) = 1/4 this is assuming the two events are independent

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (yanasidlinskiy):

Yay!!!!!!

OpenStudy (yanasidlinskiy):

Thanx:) I really need help with my next question....I will post it up in a few seconds.